OpenStudy (anonymous):

Let f(x) = -3x - 6 and g(x) = 5x + 2 Find f(x) + g(x) Choices A.) 2x-4 B.) -8x - 8 C.) -8x - 4 D.) 2x - 8

OpenStudy (anonymous):

f(x) + g(x) = (-3x-6) + (5x+2) = -3x+5x+2-6 =2x-4 ------------------>correct choice

OpenStudy (anonymous):

Thank you :)